Complete the given ordered pairs: x+y=4 (0, ) (3, ) (-2, )

Answer by macston(5194)

Complete the given ordered pairs: x+y=4 (0, ) (3, ) (-2, )
x+y=4
0+y=4
y=4
ANSWER 1: (0,4)
x+y=4
3+y=4
y=1
ANSWER 2: (3,1)
x+y=4
-2+y=4
y=6
ANSWER 3: (-2,6)